OpenStudy (anonymous):

Find a polynomial function that has the given zeros: (2,-6)

OpenStudy (anonymous):

So, you want a polynomial such that when x=2,-6 y =0

OpenStudy (anonymous):

The easiest way to do this is the following general method: if f(a) = 0, then the polynomial must contain some factor (x-a)

OpenStudy (anonymous):

That may be a little abstract, so let's do the example

OpenStudy (anonymous):

so, f(2) and f(-6) equal 0

OpenStudy (anonymous):

therefore a possible polynomial could be: \[(x-2)(x+6)\]

OpenStudy (anonymous):

as you can see, when you substitute in 2 or -6 for x, one of the 2 terms will become 0

OpenStudy (anonymous):

Now let's foil out the polynomial

OpenStudy (anonymous):

\[x^2+4x-12=y\]

OpenStudy (anonymous):

Note that this is only one possible polynomial, we could have had: \[2(x-2)(x+6)\] which would still be 0 at 2 and -6

OpenStudy (anonymous):

however, this would be equal to: \[2x^2+8x-24 = y\]

OpenStudy (anonymous):

which is also a valid solution

OpenStudy (anonymous):

So when we put it in a polynomial function it has to be the opposite what was given?

OpenStudy (anonymous):

What about (0, 12) ?